OpenStudy (anonymous):

find a quadratic function given by f(x)=a(x-h)^2+k that models the data 1940,0.25,1968,1.60,1997,5.15 need help solving for a

OpenStudy (amistre64):

ack!! not this one again..I got a program for this ;)

OpenStudy (anonymous):

lol

OpenStudy (amistre64):

i programmed this just for cases like this

OpenStudy (anonymous):

can you help me solve for a

OpenStudy (amistre64):

yep

OpenStudy (anonymous):

@amistre64 :D

OpenStudy (amistre64):

0.001301745743669519x^2 -5.0390080805461945x + 4876.675395385017

OpenStudy (amistre64):

:)

OpenStudy (amistre64):

a = something like 64.24/48957 or some rediculaous monster of a fraction

OpenStudy (amistre64):

you wanna know the min wage of 1976 right?

OpenStudy (anonymous):

so when i re write the equation it will look like f(x)= a(x-1940)^2+0.25

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

screenshot

OpenStudy (amistre64):

1976 = 2.361

OpenStudy (amistre64):

you can write the equation to be a(x-40)^2 to make the numbers easier yes

OpenStudy (anonymous):

so which of the numbers would i put in for a being that i have to show steps

OpenStudy (amistre64):

if we adjust in my program for the -40 we get this: 0.0013017457436695185(x-40)^2 + 0.011765404891539196(x-40) + 0.25

OpenStudy (amistre64):

ive got a post about this same thing a couple of days ago that I typed it all out :)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

I'd use y=ax^2+bx+c. They give you values for x and y. Like: 0.25=a(1940)^2+b(1940)+c If you use the three coordinates like that. You will have a system of equations in three variables. Solve for a, b and c. Then complete the square to get it back in vertex form which will probably be a pain.

OpenStudy (amistre64):

id have to redo it all, its a monster tho....here goes :)

OpenStudy (amistre64):

a(1940-1940)^2 + b(1940-1940) + c = .25 from this we see that c = .25 right?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

im gonna write it as the lesser numbers instead of the years ok? just be easier...

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

the other 2 equations then are: a(28)^2 +b(28) + .25 = 1.60 a(57)^2 +b(57) + .25 = 5.15

OpenStudy (amistre64):

solve for a I suppose is as likely as any other...

OpenStudy (amistre64):

a = 1.60 - .25 -b(28) --------------- now use this value in the other one (28)^2

OpenStudy (amistre64):

write down that value for a, your gonna need it :)

OpenStudy (amistre64):

a(57^2) + b(57) = 5.15 - .25 (57^2)(1.60 - .25 -b(28) ----------------------- + b(57) = 4.60 (28)^2

OpenStudy (amistre64):

(57^2)(1.35) -b(57^2)(28) + b(57)(28^2) = 4.60(28^2)

OpenStudy (amistre64):

b[-(57^2)(28) + (57)(28^2)] = 4.60(28^2) - 57^2(1.35)

OpenStudy (amistre64):

4.60(28^2) - 57^2(1.35) b = ----------------------; thats b lol (57)(28^2) - (57^2)(28)

OpenStudy (anonymous):

wow

OpenStudy (amistre64):

a = 1.60 - .25 -b(28) --------------- ; now plug the value of b into this spot (28)^2

OpenStudy (amistre64):

and thats a

OpenStudy (amistre64):

3606.40 - 4386.15 ----------------- = b 44688 - 90972

OpenStudy (amistre64):

b = 0.016847074583009247256071212514044 if I kept track of it... compare that to the one I posted earlier

OpenStudy (amistre64):

+ 0.011765404891539196; this ones better, I probably missed something earlier....

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

but if you must show your work, its just a matter of grunging your way thru it..

OpenStudy (amistre64):

The grunge work is this; formula i worked out ((p1-p2)*((n3*n3)-(n1*n1))+(p3-p1)*((n2*n2)-(n1*n1))) -------------------------------------------------- (((n3*n3)-(n1*n1))*(n1-n2)+((n2*n2)-(n1*n1))*(n3-n1))

OpenStudy (amistre64):

n1 = year1 ; p1 = pay1 and the rest is after this manner

OpenStudy (anonymous):

Lol. Parentheses fun. I always like seeing things solved for variables though.

OpenStudy (amistre64):

\[\frac{[(p1-p2)]*[(n3^2)-(n1^2)]+[(p3-p1)]*[(n2^2)-(n1^2)]}{[(n3^2)-(n1^2)]*[(n1-n2)]+[(n2^2)-(n1^2)]*[(n3-n1)]}\]

OpenStudy (amistre64):

thats 'b' lol

OpenStudy (amistre64):

you also wanna know when it was near $1 right?

OpenStudy (amistre64):

1959 is under a 1 and 1960 is a little over 1

OpenStudy (amistre64):

any questions yet? ;)

OpenStudy (amistre64):

(p2-p1) (n2-n1)*b a = ------------ - ------------- (n2^2)-(n1^2) (n2^2)-(n1^2)

OpenStudy (amistre64):

(p2-p1) - (n2-n1)*b a = ---------------------- is just as good (n2^2)-(n1^2)

OpenStudy (anonymous):

okay so this was the formula that i used to determine when minimum wage was 1.00 is this correct y=1.00(1964-1940)^2+.25 y=1.00*234^2+.25 y=(1.00*.576)+.25 y=.58+.25=1.08

OpenStudy (amistre64):

no, thats not correct minwage = y = 1 $1 is not a constant for the equation; it is a solution to it

OpenStudy (amistre64):

and you are missing your middle term

OpenStudy (amistre64):

you would take $1 from c(.25) setting the whole equation = 0; then use the quadratic formula to calculate roots as solutions and throw out the obvious wrong one

OpenStudy (anonymous):

so could you show me how to write the equation and i try to solve it again

OpenStudy (amistre64):

or trial and error it like I did ;)

OpenStudy (amistre64):

lets do this the generic way to get you to the formula I alrady created; and then all you need to do is fill in the plugs.... ok? thatll save a whole lot of little numbers getting orphaned :)

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

y = year and p = price; makes sense? a(y1)^2 + b(y1) + c = (p1)

OpenStudy (amistre64):

a(y1)^2 + b(y1) + c = (p1) Eq1 a(y2)^2 + b(y2) + c = (p2) Eq2 a(y3)^2 + b(y3) + c = (p3) Eq3

OpenStudy (amistre64):

eq1; c = (p1) -b(y1) - a(y1)^2

OpenStudy (amistre64):

laptop cord got unplugged lol

OpenStudy (amistre64):

Eq ; a(y2)^2 + b(y2) + c = (p2) a(y2)^2 + b(y2) + (p1) -b(y1) - a(y1)^2 = (p2) [a(y2)^2 - a(y1)^2] + [b(y2) -b(y1)] = (p2) - (p1) a[(y2)^2 -(y1)^2] + b[(y2) -(y1)] = (p2) - (p1) a = (p2) - (p1) - [(y2) -(y1)]b --------------------- [(y2)^2 -(y1)^2] thats what a equals a[(y2)^2 -(y1)^2] = (p2) - (p1) - [(y2) -(y1)]b

OpenStudy (amistre64):

forget that last part, its seems a bit spurious.... forgot to delete it lol

OpenStudy (amistre64):

From Eq3 we substitute all the values if for a and c a(y3)^2 + b(y3) + c = (p3) (y3^2) [(p2) - (p1) - [(y2) -(y1)]b ---------------------------- [(y2)^2 -(y1)^2] + b(y3) + (p1) -b(y1) - [(y1)^2](y3^2) [(p2) - (p1) - [(y2) -(y1)]b --------------------------------- [(y2)^2 -(y1)^2]

OpenStudy (amistre64):

we end up solving this for b and the plugging in the values that I already postd for B above :)

OpenStudy (amistre64):

you feel up to all that?

OpenStudy (anonymous):

have to be up to it lol

OpenStudy (amistre64):

I posted all the formulas; start with b go to a and then...well c is .25 for this one lol

OpenStudy (amistre64):

B = (.25-1.60) [(57^2)-(28^2)]+(5.15-.25)[(28^2)-(0^2)] ----------------------------------------------- (0 -28)[(57^2)-(28^2)]+(57-28)[(28^2)-(0^2)]

OpenStudy (anonymous):

ok this is how i did it lol(b) I believe that the quadratic function that models the data is y = a (y - 1940)2 + .25 because it seems right. (c) Estimate the minimum wage in 1976 and compare it to the actual value of $2.30. y = $1.99 (1976 – 1940)2 + .25 y = $1.99 · 362 + .25 y = ($1.99 · 1296) + .25 y = $2.58 + .25 = $2.83 Estimated $2.83 (d) Estimate when the minimum wage was $1.00. y = $1.00 (1964 – 1940)2 + .25 y = $1.00 · 242 + .25 y = ($1.00 · 576) + .25 y = $.58 + .25 = $1.08 The year 1964 (e) If current trends continue, predict the minimum wage in 2009. Compare it to the projected value of $7.25. y = $7.25 (2009 – 1940)2 + .25 y = $7.25 · 692 + .25 y = ($7.25 · 4761) + .25 y = $3.45 + .25 = $3.70 Minimum wage $3.70 @ a $3.55 difference

OpenStudy (anonymous):

i might have jacked it up sorry

OpenStudy (amistre64):

If it works for you use it ;) but dont use y for year and y for a dependant variable, its confusing lol

OpenStudy (amistre64):

the way you solve for it depends on the degree of accuracy you are looking for.

OpenStudy (anonymous):

so take y out? do i do them correct

OpenStudy (anonymous):

did i do them correct

OpenStudy (amistre64):

I dont know how you did them ;) but you shouldnt use y to mean year AND y to mean value of the function....

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

my way gives me 7.259 for min wage in 2007

OpenStudy (amistre64):

compared to 7.25 thats pretty accurate

OpenStudy (amistre64):

1960 gives me 1.006

OpenStudy (anonymous):

okay so i re did equation F (0) = .25 (0 -1.60)^2 + 5.15 .25(-1.60)^2 + 5.15 .25(2.56) + 5.15 .64 +5.15

OpenStudy (amistre64):

i dont understand the process of how you get your equation; im sure it has to do with homogeneous difference equattions tho...

OpenStudy (amistre64):

the only way I know how to solce for quadratics with three data points is to plug them into the quad equation and solve for a b and c

OpenStudy (amistre64):

and that aint what you did there :)

OpenStudy (anonymous):

okay lets scratch last equation lol

OpenStudy (amistre64):

do you know how to solve for a system of equations?

OpenStudy (amistre64):

c= .25 a(28^2) + b(28) + c = 1.60 a(58^2) + b(58) + c = 5.15 solve by substitution, elimination or matrix....

OpenStudy (amistre64):

a(28^2) + b(28) = 1.60-.25 = 1.35 a(58^2) + b(58) = 5.15-.25 = 4.90 a(28^2) + b(28) = 1.35 a = [1.35 - b(28)]/(28^2) ...................................... a(58^2) + b(58) = 4.90 ; plug in value for a

OpenStudy (amistre64):

(58^2)[1.35 - b(28)] + b(58)(28^2) = 4.90(28^2) 58^2(1.35) -b(58^2)(28) + b(58)(28^2) = 4.90(28^2) b[(-58^2)(28) +(58)(28^2)] = 4.90(28^2) - 58^2(1.35) b = 4.90(28^2) - 58^2(1.35) --------------------- (-58^2)(28) +(58)(28^2)

OpenStudy (amistre64):

plug those numbers into a calculator to get the solid value for 'b' and use it in the equation above it to solve for 'a'

OpenStudy (amistre64):

((4.90 * (28^2)) - ((58^2) * 1.35)) / (((-(58^2)) * 28) + (58 * (28^2))) = 0.014363711 Google says b = .014363711

OpenStudy (amistre64):

a = [1.35 - b(28)]/(28^2) (1.35 - (.014363711 * 28)) / (28^2) = 0.0012089491